# Theoretical Benefit and Limitation of Diffusion Language Model

### v1

@@ Abstract
Diffusion language models have emerged as a promising approach for text generation. One would naturally expect this method to be an efficient replacement for autoregressive models since multiple tokens can be sampled in parallel during each diffusion step. However, its efficiency-accuracy trade-off is not yet well understood. In this paper, we present a rigorous theoretical analysis of a widely used type of diffusion language model, the Masked Diffusion Model (MDM), and find that its effectiveness heavily depends on the target evaluation metric. Under mild conditions, we prove that when using perplexity as the metric, MDMs can achieve near-optimal perplexity in sampling steps regardless of sequence length, demonstrating that efficiency can be achieved without sacrificing performance. However, when using the sequence error rate–which is important for understanding the “correctness” of a sequence, such as a reasoning chain–we show that the required sampling steps must scale linearly with sequence length to obtain “correct” sequences, thereby eliminating MDM’s efficiency advantage over autoregressive models. Our analysis establishes the first theoretical foundation for understanding the benefits and limitations of MDMs. All theoretical findings are supported by empirical studies.
Machine Learning, ICML
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm in generative modeling, establishing state-of-the-art performance in image synthesis (Karras et al., 2022; Song et al., 2021a). Their extension to discrete domains has opened new possibilities for generating sequences, such as natural language (Campbell et al., 2022; Dieleman et al., 2022; Zheng et al., 2023; Lou et al., 2024; Campbell et al., 2024; Lovelace et al., 2024) and biological sequences (Rastogi et al., 2022; Vignac et al., 2022; Sun & Yang, 2023; Avdeyev et al., 2023). Among various discrete diffusion architectures, masked diffusion models (MDMs) (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2024)—which generate sequences by iteratively converting masks to tokens—have emerged as a prominent approach and demonstrated competitive performance across diverse language modeling tasks.
While auto-regressive models generate sequences token-by-token, discrete diffusion models can generate multiple tokens simultaneously during each step (reverse process). Therefore, it is natural to hypothesize that this parallel sampling improves generation efficiency. However, we argue that reaching a conclusion requires considering both computational cost and generation quality. Specifically, we pose the following question: do discrete diffusion models achieve superior efficiency when the generated content meets an acceptable quality standard? There may be multiple answers to this question. If diffusion models require fewer neural network executions while maintaining quality, they can offer better acceleration. Conversely, if their execution count is comparable to or exceeds that of auto-regressive models, diffusion language models may not be a better choice.
@@
To answer the above question, we leverage two complementary metrics to evaluate the efficiency of MDMs in language modeling. The first metric, token error rate (TER), quantifies token-level accuracy, which correlates with the fluency of the generated text. In practice, perplexity is a widely used metric for measuring token-level errors of language models (Jelinek et al., 1977; Devlin et al., 2019); thus, we define the metric of TER by perplexity in this paper. The second metric, sequence error rate (SER), evaluates the correctness of an entire sequence, which is crucial for reasoning tasks requiring logically correct sequences. We provide a natural definition of SER that reflects the correctness of the whole sequence. Together, these metrics enable a comprehensive evaluation of the efficiency of MDMs under both token-level and sequence-level metrics. We first provide a positive theoretical result regarding TER. We prove that under mild conditions, MDMs can achieve near-optimal TER with sampling steps regardless of the sequence length $L$. Compared to the auto-regressive model, which must be executed $L$ times to generate the sequence, MDMs demonstrate substantial efficiency gains, especially when the generation length is long.
@@
However, we show that this efficiency advantage diminishes when SER is considered. We theoretically prove that to achieve a low SER, the number of required sampling steps for MDMs scales at least linearly with sequence length. Intuitively, this limitation arises from the fact that SER, as a metric for the entire sequence, requires the generated sequence to be free of any error in the whole sequence, which forces MDMs to sample only a small number of tokens per step to mitigate such inconsistencies. As a result, the number of required sampling steps can be significant. It is notable that each MDM sampling step usually incurs a higher computational cost than an auto-regressive step under the same architecture, thus MDMs offer no efficiency advantage under this metric.
Finally, we validate our theoretical findings through comprehensive experiments. Our experiments examine MDMs trained on formal languages, including $n$-gram languages and Hidden Markov Models (HMMs), systematically analyzing the relationship between performance and efficiency under both TER and SER metrics. Additional experiments on natural language tasks, including TER evaluation on text generation and SER assessment on the GSM8k dataset (Cobbe et al., 2021), corroborate our theoretical predictions: while achieving low SER necessitates substantial sampling steps, relatively few steps suffice for TER. These results provide practical guidance for deploying diffusion language models across different applications.
@@ 2 Related Work
Discrete Diffusion Models. The auto-regressive paradigm has achieved significant success in language modeling (Dai, 2019; Floridi & Chiriatti, 2020; Achiam et al., 2023). However, its left-to-right, token-by-token generation approach is not without limitations. Notably, it faces challenges such as restricted controllability (Zhang et al., 2023) and inefficiencies in inference speed (Leviathan et al., 2023). To overcome these drawbacks, inspired by the success of diffusion models in image generation (Sohl-Dickstein et al., 2015; Song et al., 2021a; Karras et al., 2022) researchers have adapted these techniques for NLP tasks (Austin et al., 2021; He et al., 2022; Chen et al., 2022; Meng et al., 2022; Ye et al., 2023; Gulrajani & Hashimoto, 2023; Zhang et al., 2024). Discrete diffusion models, in particular, have shown promising results, achieving comparable performance with auto-regressive models across a range of NLP benchmarks.
@@
Masked Diffusion Models. Among the two primary classes of discrete diffusion models, MDMs have consistently demonstrated superior performance and scalability (Lou et al., 2024; Campbell et al., 2024). For instance, in Lou et al. (2024), the masked variant of SEDD significantly outperforms its uniform counterpart across a range of benchmarks. Similarly, Campbell et al. (2024) reports that the masked variant achieves better results in most language tasks. Furthermore, recent advancements have successfully scaled MDMs to over 1 billion parameters (Gat et al., 2024; Nie et al., 2024; Gong et al., 2024; Shi et al., 2024), underscoring their robustness and adaptability to large-scale NLP models. In this paper, we focus on MDMs, and our theoretical contributions can be applied to all MDMs, including the masked variant of discrete flow matching.
Various Metrics in NLP Tasks. Evaluation metrics in NLP tasks are inherently tied to the specific objectives and requirements of their respective domains. For general language modeling tasks, perplexity (Jelinek et al., 1977; Devlin et al., 2019) remains the metric of choice due to its ability to capture a model’s predictive performance effectively. However, domain-specific tasks often demand more specialized evaluation criteria. For instance, in machine translation (Bahdanau, 2014; Wu et al., 2016), the BLEU score is widely regarded as a standard measure of translation quality (Papineni et al., 2002), while text generation tasks (Sutskever, 2014) frequently rely on metrics such as ROUGE to assess output fidelity (Lin, 2004). Similarly, tasks requiring reasoning (Wei et al., 2022b), such as mathematics (Bubeck et al., 2023) or code generation (Roziere et al., 2023; Ouyang et al., 2023), commonly adopt accuracy as an intuitive and straightforward measure of success.
@@ 3 Masked Diffusion Language Model
3.1 Forward Process
Given a sequence ${\bm{x}}_{0}$ and a masking schedule $\alpha_{t}$, the distribution of the sequence ${\bm{x}}_{t}$ at time $t\in[0,1]$ is expressed as:
@@ 3.1 Forward Process
3.2 Reverse Process
The reverse process reconstructs a sequence from a masked version by reversing the forward dynamics. Given the sequence at time $t$ and the original sequence ${\bm{x}}_{0}$, the conditional distribution of the sequence at time $s<t$, is defined as:
@@ 3.2 Reverse Process
Here, each token is predicted independently using $p_{\theta}(x_{0}^{i}\mid{\bm{x}}_{t})$, allowing for efficient parallel sampling. However, this factorized approach imposes a significant limitation: it disregards interdependencies between tokens within the sequence. As a result, the factorized model $p_{\theta}({\bm{x}}_{0}\mid{\bm{x}}_{t})$ cannot exactly match the true reverse distribution $q({\bm{x}}_{0}\mid{\bm{x}}_{t})$ (Xu et al., 2024). In this work, we analyze the conditions under which this sampling method achieves a favorable balance between efficiency and the quality of the generated sequences.
@@ 4 Theoretical Analysis
In image generation, the primary goal is typically to produce visually appealing and seamless images (Heusel et al., 2017). Language generation is more task-specific. Depending on the application, the users may prefer fluent outputs, as in article writing, or precise and accurate reasonings, as in problem-solving tasks. In this section, we explore the sampling efficiency of MDMs in addressing various language tasks with respect to different evaluation metrics.
@@ 4.1 Notations and Problem Setting
Our investigation employs the hidden Markov model (HMM) framework to analyze natural language generation. This section establishes the formal notation and problem setting that underlies our subsequent analysis.
@@ 4.2 MDMs Can Generate Low-TER Sentences Efficiently
The proof of this theorem is presented in Appendix B.
Theorem 4.2 demonstrates that MDMs can efficiently generate sentences with high fidelity. It is notable that for a given data distribution $q$, the TER of a language model $p$ achieves its global minimum when $p=q$. To ensure a gap of at most $\epsilon$ with the optimal TER during sampling, the number of required sampling steps is bounded by $O\big{(}\frac{n-1}{\epsilon^{n}}\big{)}$.
@@ 4.3 MDMs Cannot Generate Low-SER Sentences with A Low Cost
In this subsection, we examine the SER of sampling in MDMs and highlight a fundamental limitation of MDMs in generating logically rigorous language. We begin by establishing that, with sufficient sampling steps, the MDMs have the capability to approximate a target HMM model with perfect SER.
@@
$\operatorname{SER}(p)\leq\delta.$
The complete proof of Theorem 4.3 is detailed in Section C.1. While this result establishes the theoretical capability of MDMs to achieve low SER, we still need to estimate the computational cost to achieve it. The following theorem provides a negative result for this problem.
@@
$\operatorname{SER}(p)>\frac{1}{2}.$
The proof is presented in Section C.2.
Theorem 4.4 shows that to generate sequences with low SER, the number of sampling steps in MDMs must scale at least linearly with the sequence length $L$, indicating that the number of neural network executions is comparable between MDMs and autoregressive models. However, this scaling law of MDMs typically leads to much higher computational costs compared to autoregressive models. For instance, in the case of Transformer-based architectures, each execution step in MDMs involves a quadratic computational complexity in terms of $L$, as opposed to the linear complexity of auto-regressive Transformer models in each generation step (through reusing the stored KV caches). Consequently, in accuracy-critical applications, MDMs offer no computational efficiency advantage over auto-regressive models.
Furthermore, some prior works (Sahoo et al., 2024; Ou et al., 2024) have proposed efficient sampling strategies that reuse cached outputs without requiring additional forward passes through the network when no token is modified from [m] at a given step. Nevertheless, our theoretical results remain applicable to these sampling strategies, as discussed in Appendix D.
Do TER and SER Conflict? The above results reveal that MDMs can efficiently generate low-TER sentences but may incur higher costs when evaluating the generation under SER. One might think these results are contradictory. Note that several previous works have already shown that TER (a.k.a perplexity) may not reflect a model’s true performance in solving several long-sequence understanding tasks (Huang et al., 2022; Hu et al., 2024; Luden et al., 2024). Thus, it is natural to arrive at different conclusions depending on the metric used.
Moreover, many practical scenarios have shown that the choice of evaluation metric significantly influences the conclusion of other problems. For instance, while the community has previously focused on the emergence phenomenon, recent works by Wei et al. (2022a) and Schaeffer et al. (2024) demonstrate that this phenomenon may stem from the use of non-smooth evaluation metrics. Our work further reveals that conclusions regarding the efficiency of MDMs depend heavily on the evaluation metric employed. Specifically, MDMs excel in applications where fluency is prioritized. In contrast, for reasoning-intensive tasks that demand highly accurate trajectories, MDMs may fail to offer a significant efficiency advantage over auto-regressive models.

### v2

@@ Abstract
Diffusion language models have emerged as a new approach for text generation. By enabling the parallel sampling of multiple tokens in each diffusion step, they appear to offer a more efficient alternative to auto-regressive models. However, our observations show that current open-sourced diffusion language models require more sampling steps to achieve comparable accuracy on representative tasks–resulting in even higher inference costs than their auto-regressive counterparts. To investigate whether this is an inherent limitation, we conduct a rigorous theoretical analysis of a widely adopted variant: the Masked Diffusion Model (MDM). Surprisingly, our analysis reveals that the conclusion is highly sensitive to the choice of evaluation metric. Under mild conditions, we prove that when the target is near-optimal perplexity, MDMs can achieve this goal in a constant number of sampling steps, independent of sequence length. This result demonstrates that efficiency can, in principle, be attained without compromising generation quality. However, when targeting low sequence error rate–which is important for assessing the “correctness" of a generated sequence, such as a reasoning chain–we show that in the worst case, the required sampling steps must scale linearly with sequence length, thereby eliminating the efficiency advantage. Our analysis establishes the first theoretical foundation for understanding the comparative strengths and limitations of MDMs, offering practical guidance on when to favor MDMs over auto-regressive models and vice versa.
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Song et al., 2021b) have emerged as a powerful paradigm in generative modeling, establishing state-of-the-art performance in image synthesis (Karras et al., 2022; Song et al., 2021a). Their extension to discrete domains has opened new possibilities for generating sequences, such as natural language (Campbell et al., 2022; Dieleman et al., 2022; Zheng et al., 2023; Lou et al., 2024; Campbell et al., 2024; Lovelace et al., 2024) and biological sequences (Rastogi et al., 2022; Vignac et al., 2022; Sun & Yang, 2023; Avdeyev et al., 2023). Among various discrete diffusion architectures, masked diffusion models (MDMs) (Shi et al., 2024; Sahoo et al., 2024; Ou et al., 2024)—which generate sequences by iteratively converting masks to tokens—have demonstrated competitive performance across language generation tasks.
While auto-regressive models generate sequences token-by-token, discrete diffusion models can generate multiple tokens simultaneously during each step, offering the potential for greater efficiency. However, efficiency and quality are often two sides of the same coin, and the key lies in identifying the trade-off points of different approaches. Unfortunately, as shown in Figure 1, we observed that for two recent open-sourced large MDMs, achieving performance comparable to that of left-to-right generative models incurs higher computational costs, as more sampling steps are required. This leads us to ask: do discrete diffusion models really offer a better trade-off than auto-regressive models, achieving superior efficiency while maintaining high-quality generated content? The answer may vary. If MDMs require fewer steps (i.e., neural network executions) without compromising quality, they could offer a more favorable trade-off than auto-regressive models. However, if the number of executions needed to maintain quality is similar to or exceeds that of auto-regressive models, then MDMs may not present a clear advantage.
@@
To address the above question, we present the first theoretical analysis of the efficiency of Masked Diffusion Models (MDMs) and find that the efficiency-accuracy trade-off is highly sensitive to the choice of evaluation metric. We adopt two complementary metrics to assess the efficiency of MDMs in language modeling. The first metric, token error rate (TER), quantifies token-level accuracy, which correlates with the fluency of the generated text. In practice, perplexity is a widely used metric for measuring token-level errors of language models (Jelinek et al., 1977; Devlin et al., 2019); thus, we define the metric of TER by perplexity in this paper. The second metric, sequence error rate (SER), evaluates the correctness of an entire sequence, which is crucial for reasoning tasks requiring logically correct sequences. We provide a natural definition of SER that reflects the correctness of the whole sequence. Together, these metrics enable a comprehensive evaluation of MDMs at both the token and sequence levels.
@@
We first provide a positive theoretical result regarding TER. We prove that under mild conditions, MDMs can achieve near-optimal TER with sampling steps regardless of the sequence length $L$. Compared to the auto-regressive model, which must be executed $L$ times to generate the sequence, MDMs demonstrate substantial efficiency gains, especially when the generation length is long. However, we show that this efficiency advantage diminishes when SER is considered. We theoretically prove that to achieve a low SER, in the worst case, the number of required sampling steps for MDMs must scale at least linearly with sequence length. Intuitively, this limitation arises from the fact that SER, as a metric for the entire sequence, requires the generated sequence to be free of any error in the whole sequence, which forces MDMs to sample only a small number of tokens per step to mitigate such inconsistencies. As a result, the number of required sampling steps can be significant. It is notable that each MDM sampling step usually incurs a higher computational cost than an auto-regressive step under the same architecture, thus MDMs offer no efficiency advantage.
To fully validate our theoretical findings, we conduct synthetic experiments and examine MDMs trained on formal languages, including $n$-gram languages and Hidden Markov Models (HMMs), systematically analyzing the relationship between performance and efficiency under both TER and SER metrics. All empirical results align with our theoretical predictions: For MDMs, achieving a low SER requires a significant number of sampling steps, and this requirement increases as sequence length grows. In contrast, obtaining a satisfactory TER demands fewer sampling steps, with this number remaining relatively constant regardless of sequence length. These findings offer practical guidance for selecting when to deploy diffusion language models based on specific application needs and requirements.
@@ 2 Related Work
Discrete Diffusion Models. The auto-regressive paradigm has achieved significant success in language modeling (Dai, 2019; Floridi & Chiriatti, 2020; Achiam et al., 2023). However, its left-to-right, token-by-token generation approach is not without limitations. Notably, it faces challenges such as restricted controllability (Zhang et al., 2023) and inefficiencies in inference speed (Leviathan et al., 2023). To overcome these drawbacks, inspired by the success of diffusion models in image generation (Sohl-Dickstein et al., 2015; Song et al., 2021a; Karras et al., 2022) researchers have adapted these techniques for NLP tasks (Austin et al., 2021a; He et al., 2022; Chen et al., 2022; Meng et al., 2022; Ye et al., 2023; Gulrajani & Hashimoto, 2023; Zhang et al., 2024). Discrete diffusion models, in particular, have shown promising results, achieving comparable performance with auto-regressive models across a range of NLP benchmarks.
@@
Masked Diffusion Models. Among the two primary classes of discrete diffusion models, MDMs have consistently demonstrated superior performance and scalability (Lou et al., 2024; Campbell et al., 2024). For instance, in Lou et al. (2024), the masked variant of SEDD significantly outperforms its uniform counterpart across a range of benchmarks. Similarly, Campbell et al. (2024) reports that the masked variant achieves better results in most language tasks. Based on MDMs, some sampling strategies have been proposed to enhance efficiency or generation quality (Sahoo et al., 2024; Ou et al., 2024; Wang et al., 2025; Kim et al., 2025). Furthermore, recent advancements have successfully scaled MDMs to 8 billion parameters (Gat et al., 2024; Nie et al., 2024; Gong et al., 2024; Shi et al., 2024; Nie et al., 2025; Ye et al., 2025), underscoring their robustness and adaptability to large-scale NLP models. In this paper, we focus on MDMs, and our theoretical contributions can be applied to all MDMs, including the masked variant of discrete flow matching. Notably, concurrent with our theoretical analysis of MDMs, (Li et al., 2024) also conduct an in-depth theoretical and empirical study on another class of MDMs. While their work primarily investigates the statistical complexity of learning these models, our analysis concentrates on the efficiency-accuracy trade-off of MDMs during inference.
@@ 3 Masked Diffusion Language Model
Forward process. Given a sequence ${\bm{x}}_{0}$ and a masking schedule $\alpha_{t}$, the distribution of the sequence ${\bm{x}}_{t}$ at time $t\in[0,1]$ is expressed as:
@@
Reverse process. The reverse process reconstructs a sequence from a masked version by reversing the forward dynamics. Given the sequence at time $t$ and the original sequence ${\bm{x}}_{0}$, the conditional distribution of the sequence at time $s<t$, is defined as:
@@
Here, each token is predicted independently using $p_{\theta}(x_{0}^{i}\mid{\bm{x}}_{t})$, allowing for efficient parallel sampling. However, this factorized approach imposes a significant limitation: it disregards interdependencies between tokens within the sequence. As a result, the factorized model $p_{\theta}({\bm{x}}_{0}\mid{\bm{x}}_{t})$ cannot exactly match the true reverse distribution $q({\bm{x}}_{0}\mid{\bm{x}}_{t})$ (Xu et al., 2024). In this work, we analyze the conditions under which this sampling method achieves a favorable balance between efficiency and the quality of the generated sequences. Motivated by these observations, this paper begins by re-examining the theoretical upper bounds on the performance and efficiency of MDMs.
Figure 1: Efficiency and accuracy of MDMs on GSM8K (8-shot): The left figure uses Qwen2.5-7B as the evaluation baseline and plots the relative accuracy and efficiency of Dream-v0-Base-7B and LLaDA-8B-Base across different sampling steps. The origin can be considered as the baseline result. The first quadrant demonstrates models that can simultaneously achieve higher efficiency and accuracy. From the figure, we can see that the MDMs all fall into the third quadrant, indicating lower efficiency and lower performance. The right figure uses a baseline in which the MDMs are constrained to generate the left-most masked token at each step, effectively mimicking auto-regressive decoding using the same model. Still, even under this setting, none of the configurations fall into the first quadrant.
Efficiency-accuracy trade-off of open-sourced MDMs. The efficiency-accuracy trade-off is crucial for understanding the strengths of different generative models. In Lou et al. (2024); Nie et al. (2024), researchers demonstrate that MDMs offer a more favorable trade-off than auto-regressive models in terms of perplexity—-that is, the learned MDMs can achieve competitive perplexity with higher token throughput. However, to the best of our knowledge, there has been limited empirical investigation of this trade-off in realistic tasks such as reasoning.
To address this gap, we conducted experiments on the GSM8K and MBPP benchmarks (Cobbe et al., 2021; Austin et al., 2021b), using two recently released large-scale MDMs in 2025, Dream-v0-Base-7B and LLaDA-8B-Base (Ye et al., 2025; Nie et al., 2025). We directly evaluate these checkpoints using different numbers of decoding steps, and compare the accuracy and efficiency against similar-size auto-regressive baseline Qwen2.5-7B (Team, 2024). The results on GSM8K are presented in Figure 1, the results on MBPP are presented in Section A.3. Notably, across all configurations (sampling steps), the performance and efficiency of MDMs are significantly worse than that of auto-regressive models.
A key contributing factor to this gap is that the model checkpoints were trained on different private datasets, which inherently introduces bias into the comparison. To ensure fairness, we also construct a baseline by running the MDMs in an auto-regressive manner: at each diffusion step, the model is constrained to generate the left-most masked token, thereby producing the full sequence in $L$ steps. However, even compared to this naive baseline, none of the configurations fall into the first quadrant of the plot in Figure 1—-indicating that, compared to auto-regressive decoding, MDMs may not be able to simultaneously achieve superior efficiency and improved performance. Full details are given in Appendix A.
@@ 4 Theoretical Analysis
Motivated by the observations above, this paper begins by re-examining the theoretical trade-off between the accuracy and efficiency of MDMs. In image generation, the primary goal is typically to produce visually appealing and seamless images (Heusel et al., 2017). Language generation is more task-specific. Depending on the application, the users may prefer fluent outputs, as in article writing, or precise and accurate reasoning, as in problem-solving tasks. In this section, we explore the sampling efficiency of MDMs in addressing various language tasks with respect to different evaluation metrics.
@@ 4.1 Notations and Problem Setting
Our investigation employs the hidden Markov model (HMM) framework to analyze natural language generation. This section establishes the formal notation and problem setting that underlie our subsequent analysis.
@@ 4.2 MDMs Can Generate Low-TER Sentences Efficiently
The proof of this theorem is presented in Appendix D. Theorem 4.2 demonstrates that MDMs can efficiently generate sentences with high fidelity. It is notable that for a given data distribution $q$, the TER of a language model $p$ achieves its global minimum when $p=q$. To ensure a gap of at most $\epsilon$ with the optimal TER during sampling, the number of required sampling steps is bounded by $O\big{(}\frac{n-1}{\epsilon^{n}}\big{)}$.
@@ 4.3 MDMs Cannot Generate Low-SER Sentences with A Low Cost
In this subsection, we examine the SER of sampling in MDMs and highlight a fundamental limitation of MDMs in generating logically rigorous language. We begin by establishing that, with sufficient sampling steps, the MDMs are able to approximate a target HMM model with perfect SER.
@@
Let $q$ denote any HMM, and let $p_{\mathsf{\theta}}$ represent the reverse model under an arbitrary masking schedule, where $L$ is the sequence length. Let $p$ denote the distribution over sequences generated by $p_{\mathsf{\theta}}$. Under 4.1 with a learning error $\epsilon_{\text{learning}}<O(\frac{\delta}{L})$, and given a sufficient number of reverse steps, the sequence error rate $\operatorname{SER}(p)$ of the generated text satisfies $\operatorname{SER}(p)\leq\delta.$
The complete proof of Theorem 4.3 is detailed in Section E.1. While this result establishes the theoretical capability of MDMs to achieve low SER, we still need to estimate the computational cost to achieve it. The following theorem provides a negative result for this problem.
@@
There exists an HMM $q$ over a vocabulary of size $16$ that satisfies the following conditions: for any reverse model $p_{\mathsf{\theta}}$ under 4.1 with ${\epsilon}_{\mathrm{learning}}<\frac{1}{128}$, and any masking schedule $\alpha_{t}$, let $p$ denote the distribution over sequences generated by $p_{\mathsf{\theta}}$. There exists a constant $C$ such that if the number of sampling steps satisfies $N=CL$, where $L$ is the sequence length, the SER of the generated text is lower-bounded by: $\operatorname{SER}(p)>\frac{1}{2}.$
The proof of Theorem 4.4 is presented in Section E.2. Theorem 4.4 shows that to generate sequences with low SER, the number of sampling steps in MDMs must scale at least linearly with the sequence length $L$, indicating that the number of neural network executions is comparable between MDMs and autoregressive models. However, this scaling law of MDMs typically leads to much higher computational costs compared to autoregressive models. For instance, in the case of Transformer-based architectures, each execution step in MDMs involves a quadratic computational complexity in terms of $L$, as opposed to the linear complexity of auto-regressive Transformer models in each generation step (through reusing the stored KV caches). Consequently, in accuracy-critical applications, MDMs offer no computational efficiency advantage over auto-regressive models.
Extended analysis of Theorem 4.4 Since the high SER partly arises from irreducible errors introduced by parallel sampling, it is natural to ask whether this issue can be mitigated by allowing tokens to be remasked. In Section E.3, we examine the remasking strategy recently proposed by (Wang et al., 2025), and demonstrate that Theorem 4.4 continues to hold. Furthermore, some prior works (Sahoo et al., 2024; Ou et al., 2024) have proposed efficient sampling strategies that reuse cached outputs without requiring additional forward passes through the network when no token is modified from [m] at a given step. Nevertheless, our theoretical results remain applicable to these sampling strategies, which is discussed in Section B.3.
Discussion regarding theory and GSM8k result. SER is closely related to the accuracy of MDMs in solving mathematical problems, as an incorrect chain of thoughts typically leads to an erroneous answer (Wei et al., 2022b; Yu et al., 2024; Zhu et al., 2024). Our theory reveals that MDMs incur higher computational costs when evaluating generations by SER. This finding is consistent with results presented in Figure 1. Both our theoretical insights and these prior experiments demonstrate that when MDMs are applied to math problems, their efficiency-performance trade-off is less favorable than that of AR models. For further discussion, please refer to Section B.2.
Practical guideline of ARs v.s. MDMs. Based on the theoretical results, conclusions regarding when to favor MDMs depend heavily on the evaluation metric employed. Specifically, MDMs excel in applications where fluency is prioritized. In contrast, for reasoning-intensive tasks that demand highly accurate trajectories, MDMs may fail to offer a significant efficiency advantage over auto-regressive models. As a result, MDMs are better suited for fluent generation tasks, while ARs remain the preferred choice for tasks that require precise, step-by-step reasoning.